OpenStudy (anonymous):

Find e of an ellipse if the distance between its foci is same as the length of its latus-rectum?

OpenStudy (anonymous):

My answer is Case I e = 4/ (1+sqrt (17) ) Case II e = 4/ (1-sqrt (17) )

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Am i correct?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Luis_Rivera